33.3 What is the de Broglie wavelength of: (a) an electron (melectron electron 9.1 x 10-31 kg) travelling at 15 kms-1? (b) an electron with a kinetic energy of 1 eV? (c) a proton (mproton -1.67x 10-27 kg travelling at 15kms-1? (d) a proton with a kinetic energy of 1 eV? (e) an elephant (melephant 10 tonnes) travelling at 15 km h-1? (D an elephant with a kinetic energy of 1 evi
Answer #1

The de Broglie wavelength (lambda) equation is

lambda = rac{h}{p}

Where, h = 6.63 x 10-34 Js is plank's constant and

p is momentum of the particle

(a) Given,

mass of electron, me = 9.1 x 10-31 kg

velocity of electron = 15 km/s = 15000 m/s

So, momentum of electron = 15000 x  9.1 x 10-31 = 1.365 x 10-26 kg.m/s

Therefore, de Brolie wavelength of electron is,

  6.63 x 10-34 1.365 x 10-26 .85 x 10-m 48.5 nm

(b) Given, the kinetic energy (E) of electron = 1 eV = 1.602 x 10-19 J

Momentum of the electron can be calculated using this equation :

p = sqrt{2m_{e}E}

or, p = V2 x 9.1 x 10-31 x 1.602 x 10-19-V29.1564 x 10-50

5.34 x 10-25 kgm/s

So, de Broglie wavelength of electron is,

6.63 x 10-34 5.34 x 10-25 1.23 x 10- - 1.23 nm

(c)  

Given,

mass of proton, mp = 1.67 x 10-27 kg

velocity of electron = 15 km/s = 15000 m/s

So, momentum of electron = 15000 x 1.67 x 10-27 = 2.5050 x 10-23 kg.m/s

Therefore, de Brolie wavelength of electron is,

  6.63 x 10-34 2.5050 x 10-23 = 2.65 x 10-11 m 0.265 (Angstrom)

(d)  

Given, the kinetic energy (E) of proton = 1 eV = 1.602 x 10-19 J

Momentum of the proton can be calculated using this equation :

p = sqrt{2m_{p}E}

or,   p = V2 x 1.67 x 10-27 x 1.602 x 10-19-V535 x 10-46

= 2.31 imes 10^{-23} ,, kg m/s

So, de Broglie wavelength of proton is,

6.63 x 10-34 2.31 x 10-23 -2.87 x 10-11 mal. 287 Ao

(e)  

Given,

mass of elephant, melephant = 10000 kg

velocity of elephant = 15 km/hr = 4.167 m/s

So, momentum of elephant = 4.167 x 10000= 4.167 x 104 kg.m/s

Therefore, de Brolie wavelength of elephant is,

  6.63 x 10-34 4.167 x 101.59 x 10-38,m 1.59 x 10-38 m

(f)  

Given, the kinetic energy (E) of elephant = 1 eV = 1.602 x 10-19 J

Momentum of the elephant can be calculated using this equation :

2melephant

or,   p = sqrt{2 imes 10000 imes 1.602 imes 10^{-19}} = sqrt{32.04 imes 10^{-16} }

5.66 x 10-8 kgm/s

So, de Broglie wavelength of elephant in this case is,

6.63 x 10-34 5.66 x 10-8-1.17 x 10-26

For any doubt please comment and please give an up vote. Thank you.